# Theoretical advances in electroweak, Higgs, and top physics at the LHC

### v1

@@ 1 Introduction
Since its discovery in 2012, the Higgs boson has been one of the main driving force of the physics programme of the Large Hadron Collider (LHC). In particular, a lot of emphasis has been put on studying Higgs-boson properties [1]. It is therefore natural that, even after more than 10 years, Higgs physics is still playing an important role at the LHC.
This situation will be reflected in these proceedings by covering several recent state-of-the-art calculations, all related to some extent to Higgs physics at the LHC. In particular, a combined calculation for triboson and WH production will be highlighted. Recent progresses on electroweak (EW) corrections for double-Higgs production will be shortly reviewed. Approximate predictions at NNLO QCD for the total cross section of Higgs production in association with a top-antitop pair will be discussed along with progresses regarding the computation of the corresponding two-loop contribution. Finally, new results for ttW production, which is an important background in ttH measurements, will be shown.
@@ 2 Triboson and WH production
In Ref. [5], the full NLO QCD and EW corrections were computed for both the EW and QCD production of two same-sign W boson in association with two jets which is the golden channel for vector-boson scattering (VBS) at the LHC. This process being a particularly important probe of the EW sector. This very final state can actually also be obtained from triboson- and WH-production mechanisms where one W boson is decaying hadronically, as illustrated in Fig. 1.
@@
In Ref. [6], a new full calculation has been performed for a triboson phase space inspired by a recent ATLAS analysis [7]. The first interesting finding of this work is that the triboson phase space used by the ATLAS collaboration does not only enhance triboson production but also WH production at the level of $40\%$ of the fiducial cross section. This can also be seen in the differential distributions of Fig. 2 where the various production mechanisms display a non-trivial interplay across the whole phase space. Making a distinction between these different mechanisms is therefore somehow artificial.
@@
Beyond the purely phenomenological interest of the calculation, there are also more theoretical questions that motivated this study. For example, in Ref. [8], it has been shown that large EW corrections are an intrinsic feature of VBS at the LHC. On the other hand, triboson production has shown rather moderate EW corrections. Given that these two processes share the same phase final state, these two statements seem contradictory. Reference [6] has shown that while EW corrections are still large for $t$-channel contributions (like for VBS), in the triboson phase-space they are suppressed with respect to the $s$-channel ones that are enhanced but display small EW corrections.
Along with a fixed-order analysis, state-of-the-art predictions featuring parton-shower (PS) corrections using the public Monte Carlo program Sherpa [9] has been presented for both the EW and the QCD production. The parton-shower corrections are about $10\%$ for the fiducial cross section but can be significantly more pronounced at the differential level as illustrated in Fig. 3.
@@ 3 HH production
The double-Higgs production is most likely the next Standard Model processes to be measured at the LHC for the first time. It is particularly crucial for the exploration of the Electroweak Symmetry Breaking Mechanism by providing sensitivity to triple-Higgs coupling.
Given that the partonic process ${\rm g}{\rm g}\to{\rm H}{\rm H}$ is loop-induced, LO predictions are already a one-loop calculation while the next-to-leading oder (NLO) predictions feature two-loop contributions. For the EW corrections, the real radiation, which is a photon radiation, is actually vanishing due to Furry’s theorem. In addition, weak radiations (of W and Z bosons) are not needed for infrared finiteness of the cross section and are typically rejected in experimental analysis. It also means that the calculation of the NLO EW corrections only contains the two-loop virtual contribution.
This particularly challenging calculation has been tackled by two independent groups. In Ref. [10] the exact calculation was presented while in Ref. [11], an expansion in the large top-mass limit was used. Note that very recently, Ref. [12] also presented results for the EW corrections to the top-Yukawa and the self-coupling contributions.
Example contributing Feynman diagrams are shown in the left hand-side of Fig. 4. The top diagrams display the LO contributions while the others show the two-loop EW virtual contributions. On the right-hand side of Fig. 4, the LO and NLO EW predictions for the full calculation of Ref. [10] are shown. The EW corrections display a typical behaviour for the LHC, namely that the corrections grow negatively large in the high-energy limit under the influence of Sudakov logarithms. For example, the corrections are at the level of $-10\%$ at $800$ GeV.
Figure 4: Sample Feynman diagram for ${\rm g}{\rm g}\to{\rm H}{\rm H}$ at the LHC at LO and NLO EW (left). Transverse momentum of one of the two Higgs bosons at LO and NLO EW (right). All figures are taken from Ref. [10].
@@
In Ref. [13], a full computation at next-to-next-to-leading order (NNLO) QCD accuracy was presented. The only approximation made in the calculation is in the two-loop contribution, which is a particularly challenging part of this calculation. For this two-loop virtual contribution, the authors reverted to the so-called soft Higgs approximation which is valid for scales much larger than the transverse momentum of the Higgs boson. Such an approximation is appropriate for inclusive numbers and has been found to be $1\%$ accurate in this case. In Fig. 5, inclusive cross sections as a function of the centre-of-mass energy at the LHC are displayed at NNLO QCD accuracy.
Figure 5: Inclusive cross sections as a function of the centre-of-mass energy in perturbative QCD for ttH production at the LHC. The distribution is taken from Ref. [13].
@@
Figure 6: Topologies contribution to the two-loop amplitude for ttH production at the LHC (left). Example numerical results for squared amplitudes at $\mathcal{O}\!\left(\epsilon^{-1}\right)$ for NNLO QCD for the $q\bar{q}$ channel (right). All figures are taken from Ref. [14].
Finally, it worth mentioning Ref. [15] where the $N_{f}$ part of the two-loop amplitude (i.e. with closed fermion loops) of the $q\bar{q}$ channel has been computed. Some contributing Feynman diagrams are provided on the left hand-side of Fig. 7. It is important to point out that, as opposed to the two previous cases, this calculation provides exact numerical results (as illustrated for example on the right hand-side of Fig. 7). It constitutes a proof of concept for the full calculation. Note that the drawback of this approach is that the computation of each phase-space point takes several minutes. This means that in order to use such results for Monte Carlo integration, interpolation grids will be needed.
@@
This calculation is therefore for a $2\to 8$ scattering process and hence particularly challenging numerically. One interesting aspect of the NLO corrections is that it has been observed that subleading corrections which are denoted by NLO3 (QCD corrections to the EW production) are surprisingly large [18, 17] for on-shell top quarks and a W boson. Following power-counting arguments, one expects this correction to be rather suppressed. It is actually not the case due to ${\rm W}{\rm t}\to{\rm W}{\rm t}$ scatterings arising at this order.
@@
Figure 9: Differential distributions for off-shell ttW production at the LHC a full NLO accuracy. All distributions are taken from Ref. [16].
@@ 6 Conclusion
In these proceedings, I have highlighted recent computations that illustrate the current frontier of theoretical predictions at the LHC within the Standard Model. The interested reader is referred to the original work for more information. The processes mentioned above were all related to some extend to Higgs-boson physics due to its prevalence at the LHC. In detail, the list of the processes reported with their corresponding features reads
@@
These exemplary calculations constitute corner stones of the precision programme currently ongoing at the LHC. They are therefore decisive information for Standard Model tests that will be performed during the run III of the LHC and beyond.

### v3

@@ 1 Introduction
Since its discovery in 2012, the Higgs boson has been one of the main driving forces of the physics programme of the Large Hadron Collider (LHC). In particular, a lot of emphasis has been put on studying Higgs-boson properties [1]. It is therefore natural that, even after more than 10 years, Higgs physics is still playing an important role at the LHC.
This situation will be reflected in these proceedings by covering several recent state-of-the-art calculations, all related to some extent to Higgs physics at the LHC. In particular, a combined calculation for triboson and WH production will be highlighted. Recent progress on electroweak (EW) corrections for double-Higgs production will be shortly reviewed. Approximate predictions at NNLO QCD for the total cross section of Higgs production in association with a top-antitop pair will be discussed along with progresses regarding the computation of the corresponding two-loop contribution. Finally, new results for ttW production, which is an important background for ttH measurements, will be shown.
@@ 2 Triboson and WH production
In Ref. [5], full NLO QCD and EW corrections were computed for both the EW and QCD production of two same-sign W boson in association with two jets which is the golden channel for vector-boson scattering (VBS) at the LHC. This process being a particularly important probe of the EW sector. This very final state can actually also be obtained from triboson- and WH-production mechanisms where one W boson is decaying hadronically, as illustrated in Fig. 1.
@@
In Ref. [6], a new full calculation has been performed for a triboson phase space inspired by a recent ATLAS analysis [7]. The first interesting finding of this work is that the triboson phase space used by the ATLAS collaboration does not only contain triboson production but also WH production at the level of $40\%$ of the fiducial cross section. This can also be seen in the differential distributions of Fig. 2 where the various production mechanisms display a non-trivial interplay across the whole phase space. Making a distinction between these different mechanisms is therefore somehow artificial.
@@
Beyond the purely phenomenological interest of the calculation, there are also more theoretical questions that motivated this study. For example, in Ref. [8], it has been shown that large EW corrections are an intrinsic feature of VBS at the LHC. On the other hand, triboson production has shown rather moderate EW corrections. Given that these two processes share the same final state, these two statements seem contradictory. Reference [6] has shown that while EW corrections are still large for $t$-channel contributions (like for VBS), they are suppressed with respect to the $s$-channel ones that are enhanced but display small EW corrections.
Along with a fixed-order analysis, state-of-the-art predictions featuring parton-shower (PS) corrections using the public Monte Carlo program Sherpa [9] has been presented for both the EW and the QCD production as illustrated in Fig. 3. The parton-shower corrections are about $10\%$ for the fiducial cross section but can be significantly more pronounced for some differential distributions.
@@ 3 HH production
The double-Higgs production is certainly one of the next processes to be measured at the LHC. It is particularly crucial for the exploration of the Electroweak Symmetry Breaking Mechanism by providing sensitivity to triple-Higgs coupling.
Given that the partonic process ${\rm g}{\rm g}\to{\rm H}{\rm H}$ is loop-induced, LO predictions are already a one-loop calculation while the next-to-leading oder (NLO) predictions feature two-loop contributions. For the EW corrections, the real radiation, which is a photon radiation, is actually vanishing due to Furry’s theorem. Note that weak radiations (W,Z) are not needed for infrared finiteness of the cross section and are typically rejected in experimental analysis. It also means that the calculation of the NLO EW corrections only contain the two-loop virtual contribution.
This particularly challenging calculation has been tackled by several independent groups. In Ref. [10] the exact calculation was presented while in Refs. [11, 12, 13, 14, 15] several partial results and/or expanded results have been made public.
Example contributing Feynman diagrams are shown on the left hand-side of Fig. 4. The top diagrams display the LO contributions while the others show the two-loop EW virtual contributions. On the right-hand side of Fig. 4, the LO and NLO EW predictions for the full calculation of Ref. [10] are shown. The EW corrections display a typical behaviour for the LHC, namely that the corrections grow negatively large in the high-energy limit under the influence of Sudakov logarithms. For example, the corrections are at the level of $-10\%$ at $800$ GeV.
Figure 4: Sample Feynman diagram for ${\rm g}{\rm g}\to{\rm H}{\rm H}$ at the LHC at LO and NLO EW (left). Distribution in the transverse momentum of one of the two Higgs bosons (right). All figures are taken from Ref. [10].
@@
In Ref. [16], a full computation at next-to-next-to-leading order (NNLO) QCD accuracy was presented. The only approximation made in the calculation is in the two-loop contribution, which is the most challenging part of this calculation. For the two-loop virtual contribution, the authors reverted to the so-called soft Higgs approximation which is valid for scales much larger than the transverse momentum of the Higgs boson. Such an approximation is appropriate for inclusive numbers and has been found to be $1\%$ accurate in this case. In Fig. 5, inclusive cross sections as a function of the centre-of-mass energy at the LHC are displayed at NNLO QCD accuracy.
Figure 5: Inclusive cross sections as a function of the centre-of-mass energy in perturbative QCD for ttH at the LHC. The figure is taken from Ref. [16].
@@
Figure 6: Topologies contributing to the two-loop amplitude for ttH production at the LHC (left). Example numerical results for squared amplitudes at $\mathcal{O}\!\left(\epsilon^{-1}\right)$ for NNLO QCD for the $q\bar{q}$ channel (right). All figures are taken from Ref. [17].
Finally, it is worth mentioning Ref. [18] where the $N_{f}$ part of the two-loop amplitude (i.e. with closed fermion loops) of the $q\bar{q}$ channel has been computed. Some contributing Feynman diagrams are provided on the left hand-side of Fig. 7. It is important to point out that, as opposed to the two previous cases, this calculation provides exact numerical results (as illustrated for example on the right hand-side of Fig. 7). It constitutes a proof of concept for the full calculation. Note that the drawback of this approach is that the computation of each phase-space point takes several minutes. This means that in order to use such results for Monte Carlo integration, interpolation grids will be needed. Finally, note that progress towards a calculation of two-loop virtual contributions without approximation by other groups has been started [19, 20].
@@
This calculation is therefore for a $2\to 8$ process and hence particularly challenging. One interesting aspect of the NLO corrections is that it has been observed that subleading corrections which are denoted by NLO3 (QCD corrections to the EW production) are particularly large [23, 22] for on-shell top quarks and a W boson. Following power-counting arguments, one expects this correction to be rather suppressed. It is actually not the case due to ${\rm W}{\rm t}\to{\rm W}{\rm t}$ scattering arising at this order.
@@
Figure 9: Differential distributions for off-shell ttW at the LHC a full NLO accuracy. All diagrams are taken from Ref. [21].
@@ 6 Conclusion
In these proceedings, I have highlighted recent computations that illustrate the current frontier of theoretical predictions at the LHC within the Standard Model. The interested reader is referred to the original work for more information. The processes mentioned above were all related to some extent to Higgs-boson physics due to its prevalence at the LHC. In detail, the list of the processes reported with their corresponding features reads
@@
These calculations constitute corner stones of the precision programme currently ongoing at the LHC. They are therefore decisive information for Standard Model tests that will be performed during the run III of the LHC and beyond.